# Structured Initialization for Vision Transformers

### v1

@@ Appendix C Proof for Proposition 1
With Remark 1, we can suppose the rank of $\mathbf{X}\,{\approx}\,\mathbf{Z}\mathbf{A}$ is $k$, where $\mathbf{Z}\,{=}\,[\mathbf{z}_{1},\ldots,z_{k}]$ and $k\,{\ll}\,D$, as illustrated in Fig. 2. We then obtain
@@
Appendix D Convolutional Represetation Matrix
In Sec. 3, we interchangeably use the terms convolution filter $\mathbf{h}$ and convolution matrix $\mathbf{H}$. Additionally, we represent the impulse filter as a convolutional matrix. Here, we offer a detailed explanation of the relationship between the convolutional filters and the convolutional matrices.
@@ Appendix D Convolutional Represetation Matrix
Appendix E Pseudo Code for Solving $\mathbf{Q}_{\text{init}}$ and $\mathbf{K}_{\text{init}}$
1:$\mathbf{P}$triangleright Input positional encoding
2:$d,f,\alpha,\beta,\gamma$triangleright Hyperparameters
3:$\mathbf{Q}_{\text{init}},\,\mathbf{K}_{\text{init}}$triangleright Initialized attention parameters
4:$N,D\leftarrow\text{shape}(\mathbf{P})$
5:$\mathbf{H}_{\text{impulse}}\leftarrow\text{ImpulseConvMatrix}(N,f)$triangleright Build 2D impulse convolution matrix
6:$\widetilde{\mathbf{M}}\leftarrow\alpha\mathbf{H}_{\text{impulse}}+\beta\mathbf{Z}$triangleright Get ideal map
7:$\widetilde{\mathbf{X}}\leftarrow\text{LayerNorm}(\mathbf{P})$triangleright Get pseudo input
8:$\mathbf{P}_{\text{inv}}\leftarrow(\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}})^{-1}\widetilde{\mathbf{X}}$triangleright Get pseudo inverse of $\mathbf{P}$
9:$\hat{\mathbf{M}}\leftarrow\mathbf{P}_{\text{inv}}\widetilde{\mathbf{M}}\mathbf{P}_{\text{inv}}^{\top}$triangleright Get patterns $\mathbf{Q}\mathbf{K}^{T}$
10:$\mathbf{U},\mathbf{s},\mathbf{V}^{\top}\leftarrow\text{SVD}(\hat{\mathbf{M}})$triangleright SVD of $\hat{\mathbf{M}}$
11:$\widetilde{\mathbf{Q}}_{\text{init}}\leftarrow\mathbf{U}\sqrt{\mathbf{s}}$, $\widetilde{\mathbf{K}}_{\text{init}}\leftarrow\mathbf{V}\sqrt{\mathbf{s}}$
@@ Appendix E Pseudo Code for Solving $\mathbf{Q}_{\text{init}}$ and $\mathbf{K}_{\text{init}}$
13:$\widetilde{\mathbf{Q}}_{\text{init}}\leftarrow\widetilde{\mathbf{Q}}_{\text{init}}/\text{norm}(\mathbf{Q})$, $\widetilde{\mathbf{K}}_{\text{init}}\leftarrow\widetilde{\mathbf{K}}_{\text{init}}/\text{norm}(\mathbf{K})$
14:$\mathbf{Q}_{\text{init}}\leftarrow\gamma\widetilde{\mathbf{Q}}_{\text{init}}$,  $\mathbf{K}_{\text{init}}\leftarrow\gamma\widetilde{\mathbf{K}}_{\text{init}}$
@@
Appendix F Training Details and Training Curves for ViT-Base
In the main paper, we found it challenging to reproduce the ViT-Base performance reported in DeiT [26], even when strictly following their specified training settings. Upon a careful examination of code differences across various versions of the timm [31] library, we identified two critical discrepancies that likely contributed to the performance gap: (a) Learning rate scaling strategy: In DeiT, the learning rate is linearly scaled with respect to the batch size: $\text{lr}_{\text{scaled}}\,{=}\,\text{lr}_{\text{{base}}}\,{\times}\,\frac{\text{batch size}}{512}$. In contrast, the current version of timm uses a square root scaling rule as the default for the AdamW optimizer: $\text{lr}_{\text{scaled}}\,{=}\,\text{lr}_{\text{{base}}}\,{\times}\,\sqrt{\frac{\text{batch size}}{512}}$. This discrepancy in the default setting leads to different effective learning rates, even when all the other hyperparameters are identical, and can substantially affect performance. (b) Repeated data augmentation setting: DeiT emphasizes the importance of the repeated data augmentation strategy, stating a drop in top-1 accuracy from $81.8\%$ to $76.5\%$ when it is disabled. However, they did not specify the exact augmentation weighting value in their original paper. After inspecting their specific version of code, we discovered that the default value for the repeated data augmentation was set to 3, whereas we used 1 in our main experiments, which may partially explain the performance discrepancy. We also aligned several other minor settings, such as the minimum learning rate during warm-up and at the end of training, as well as an additional 10 epochs for cool-down. However, we believe that these factors have only a marginal effect on the final performance, while the two aforementioned reasons remain the primary contributors to the observed discrepancy.
Table 5: Ablation study on training settings of ViT-Base on ImageNet-1K dataset. Note that we compare the performance to the default initialization method (shaded in gray). The performance we reported in the main paper is on the first row (shaded in cyan).
Scaled LR Repeated Augment Default Mimetic Ours (impulse) 7e-4 1.0 76.25 79.12   2.87$\uparrow$ 80.48   4.23$\uparrow$ 1e-3 1.0 80.09 80.17   0.08$\uparrow$ 81.55   1.07$\uparrow$ 1e-3 3.0 81.24 80.56   0.68$\downarrow$ 81.83   0.59$\uparrow$
@@ Appendix F Training Details and Training Curves for ViT-Base
We present an ablation study on these settings in Tab. 5. With the two settings properly adjusted, we achieved a top-1 accuracy of $81.24\%$, which is comparable to the $81.8\%$ reported in DeiT [26]. Under the same settings, our impulse initialization method still achieves an accuracy of $81.83\%$ with a $0.59\%$ improvement, showcasing the robustness to different training hyperparameters, and still outperforming the mimetic initialization. Furthermore, our structured initialization method demonstrates robustness across different training configurations, consistently yielding over $80\%$ accuracy and outperforming the mimetic initialization in all cases.
In Fig. 4, we also present the training accuracy curves across epochs for different initialization methods with three different training configurations. Our impulse initialization consistently outperforms the default initialization throughout the entire training process, with a particularly significant advantage in the final training stage. While the mimetic initialization shows relatively faster initial convergence, it ultimately degrades performance in the final training stage across different training configurations.
Appendix G Additional Results for Small and Medium-Scale Datasets
G.1 Dataset Scale
For completeness, we provide the detailed dataset scales used in the main paper in Tab. 6. The ordering of the datasets follows that in [33].
@@ G.1 Dataset Scale
[…]
G.2 Additional Statistical Results
Here we provided the mean and standard deviation of 5 runs for ViT-Tiny with different initialization methods on small and medium-scale datasets in Fig. 5, which serves as additional statistical results of Tab. 1 in the main paper. Notably, our initialization method still outperforms other methods across small and medium-scale datasets. Especially for the smaller-scale datasets, our methods shows larger performance improvements, aligning with our findings in the main paper.
@@ G.2 Additional Statistical Results
Appendix H Experimental Validation of the Propositions and Corollaries in Sec. 3
In this section, we conduct a series of preliminary experiments on CIFAR-10 to validate the propositions and corollaries in Sec. 3 of the main paper, with particular focus on demonstrating that in ConvMixer, fixed random impulse spatial filters can achieve comparable performance to learned filters. All experiments adhere to the training protocol proposed in [34], which is specifically tailored for evaluating diverse architectures on small-scale datasets such as CIFAR-10.
To support our theoretical findings in Sec. 3 concerning the effectiveness of random filters, we train ConvMixer [27] models with an embedding dimension of 256, a depth of 8, and a patch size of 2 on the CIFAR-10 dataset, using spatial filter sizes of 3, 5, and 8. We also include a variant with an embedding dimension of 512 to examine the impact of feature width. We evaluate the end-to-end trained ConvMixer alongside three initialization strategies: random (Corollary 1), impulse (Corollary 2), and box (Corollary 3). Note that in all three cases, only the spatial convolution filters are initialized—the models are evaluated without any training. The box filters use all-one values, effectively performing average pooling. The results are summarized in Tab. 7.
@@
Kernel Size Embedding Dimension $=$ 256 Embedding Dimension $=$ 512 Trained Random Impulse Box Trained Random Impulse Box 3 91.76 90.72 90.68 81.70 92.82 92.15 92.20 81.90 5 92.69 90.87 90.41 80.57 93.90 92.72 91.91 81.19 8 92.34 88.12 87.82 78.95 92.96 90.09 89.61 80.10
Appendix I Attention Maps
Here we provide additional visualization of the attention maps for all 12 layers in Fig. 6. In particular, our structured initialization method offers different attention peaks on various heads, showing alignment with the impulse structures, while the mimetic initialization only presents main-diagonal peaks, and the default initialization shows little to no patterns. As stated in the main paper, both mimetic and default initialization methods use identical initialization for all attention heads.
Figure 6: Visualization of attention maps in ViT-T using our impulse initialization method, mimetic [28], and default [31] initializations.

### v2

@@ Appendix B Proof for Proposition 1
With Remark˜1, we can suppose the rank of $\mathbf{X}\,{\approx}\,\mathbf{Z}\mathbf{A}$ is $k$, where $\mathbf{Z}\,{=}\,[\mathbf{z}_{1},\ldots,z_{k}]$ and $k\,{\ll}\,D$, as illustrated in Fig.˜2. We then obtain
@@
Appendix C Convolutional Represetation Matrix
In Sec.˜3, we interchangeably use the terms convolution filter $\mathbf{h}$ and convolution matrix $\mathbf{H}$. Additionally, we represent the impulse filter as a convolutional matrix. Here, we offer a detailed explanation of the relationship between the convolutional filters and the convolutional matrices.
@@ Appendix C Convolutional Represetation Matrix
for $i\,{=}\,1,2,\,{\dots}\,,f$. $\mathbf{F_{i}}$s are circulant matrices and $\mathbf{H}$ is a block circulant matrix with circulant block (BCCB). Note that convolutions may employ various padding strategies, but the circulant structure remains consistent. Here, we show the convolution matrix without any padding as an example.
Appendix D Pseudo Code for Solving $\mathbf{Q}_{\text{init}}$ and $\mathbf{K}_{\text{init}}$
@@
[…]
@@
Appendix E Training Details and Training Curves for ViT-Base
At first, we found it challenging to reproduce the ViT-Base performance reported in DeiT [touvron2021training], even when strictly following their specified training settings. Upon a careful examination of code differences across various versions of the timm [rw2019timm] library, we identified two critical discrepancies that likely contributed to the performance gap: (a) Learning rate scaling strategy: In DeiT, the learning rate is linearly scaled with respect to the batch size: $\text{lr}_{\text{scaled}}\,{=}\,\text{lr}_{\text{{base}}}\,{\times}\,\frac{\text{batch size}}{512}$. In contrast, the current version of timm uses a square root scaling rule as the default for the AdamW optimizer: $\text{lr}_{\text{scaled}}\,{=}\,\text{lr}_{\text{{base}}}\,{\times}\,\sqrt{\frac{\text{batch size}}{512}}$. This discrepancy in the default setting leads to different effective learning rates, even when all the other hyperparameters are identical, and can substantially affect performance. (b) Repeated data augmentation setting: DeiT emphasizes the importance of the repeated data augmentation strategy, stating a drop in top-1 accuracy from $81.8\%$ to $76.5\%$ when it is disabled. However, they did not specify the exact augmentation weighting value in their original paper. After inspecting their specific version of code, we discovered that the default value for the repeated data augmentation was set to 3, whereas we used 1 in our main experiments, which may partially explain the performance discrepancy. We also aligned several other minor settings, such as the minimum learning rate during warm-up and at the end of training, as well as an additional 10 epochs for cool-down. However, we believe that these factors have only a marginal effect on the final performance, while the two aforementioned reasons remain the primary contributors to the observed discrepancy.
Table 5: Ablation study on training settings of ViT-Base on ImageNet-1K dataset. Note that we compare the performance to the default initialization method (shaded in gray).
[…]
@@ Appendix E Training Details and Training Curves for ViT-Base
We present an ablation study on these two main settings difference in Tab.˜5. In Fig.˜5, we also present the training accuracy curves across epochs for different initialization methods with three different training configurations. Our impulse initialization consistently outperforms the default initialization throughout the entire training process, with a particularly significant advantage in the final training stage. While the mimetic initialization shows relatively faster initial convergence, it ultimately degrades performance in the final training stage. Our structured initialization method demonstrates robustness across different training configurations, consistently yielding over $80\%$ accuracy in all cases.
Appendix F Additional Results for Small and Medium-Scale Datasets
F.1 Dataset Scale
For completeness, we provide the detailed dataset scales used in the main paper in Tab.˜6. The ordering of the datasets follows that in [xu2024initializing].
@@ F.1 Dataset Scale
represents large-scale datasets, $\blacktriangle$ represents medium-scale datasets, and represents small-scale datasets. The datasets are ranked based on their scales following [xu2024initializing]. A ’+{+}’ means we use both train and validation dataset in training. Note for Flowers and Pets datasets, we use both training and validation data for training.
Dataset $\downarrow$ #Classes #Train images in each class #Train images in total #Images in total $\blacksquare$  ImageNet-1K 1K $\sim$}\,1K 1.3M 1.4M $\blacktriangle$ Food-101 101 750 75K 101K $\blacktriangle$ CIFAR-10 10 5K 50K 60K $\blacktriangle$ CIFAR-100 100 500 50K 60K  STL-10 10 500 5K 13K  Flowers 102 10+10 2K 8K  Pets 37 50+50 3.7K 7.3K
F.2 Additional Statistical Results
Here we provided the mean and standard deviation of 5 runs for ViT-Tiny with different initialization methods on small and medium-scale datasets in Fig.˜6, which serves as additional statistical results of Tab.˜1 in the main paper. Notably, our initialization method still outperforms other methods across small and medium-scale datasets. Especially for the smaller-scale datasets, our methods shows larger performance improvements, aligning with our findings in the main paper.
@@ F.2 Additional Statistical Results
F.3 Model Convergence Rate
We would like to clarify that while impulse filters mimic convolutional locality, they do not reuse weights like CNNs do—which may explain the fast convergence of CNNs. Interestingly, we did observe a faster convergence when using the mimetic or impulse initialization to replace the vanilla ViT model. We give an example of the training accuracy of ViT-Tiny with different initialization methods on CIFAR-10 during the optimization stage, as shown in Tab.˜7. We can clearly see a faster convergence of mimetic and our impulse initialization methods compared to the default initialization used in vanilla ViTs. One intuitive explanation is that the mimetic initialization yields ${\sim}4{\times}$ the norm values of the default initialization, while our initialization method has ${\sim}2{\times}$ the norm values of the default initialization. Different norm values will affect the gradient magnitudes and early training dynamics.
Table 7: Comparison of performance when training ViT-Tiny on CIFAR-10 at different epochs. Our method consistently achieves higher accuracy than both default and mimetic initialization baselines.
Epoch Default Mimetic Ours (Impulse) 10 47.95 49.05 54.62 50 61.51 63.62 69.51 100 76.87 77.64 81.87
F.4 Performance Improvement Increases as the Model Size Increases.
In the main paper Tab.˜2, we also observe that as the model scale increase (i.e., the number of heads increases), our method becomes relatively more expressive. Theoretically, as explained in Sec.˜3, to replace the spatial kernels with fixed filters, the initialized attention heads must span the spatial filter basis. Therefore, for a $3{\times}3$ filter used in the experiments, this requires at least $9$ independent heads to span the kernel space. Please note that in ConvMixer, each channel uses an independent filter, which satisfies this criterion. However, ViT models share the same filters within each head, which need a separate discussion: 1) ViT-Tiny has $3$ heads, ViT-Small has $6$ heads—insufficient to span to the kernel space; 2) ViT-Base has $12$ heads—sufficient for replacing the spatial kernels. This analysis fully supports this performance improvement increase observation. However, as the motivation of the paper also suggested, training even larger ViT models on small datasets is notoriously challenging due to optimization instability and is out of the scope of this paper.
Appendix G Experimental Validation of the Propositions and Corollaries in Sec.˜3
In this section, we conduct a series of preliminary experiments on CIFAR-10 to validate the propositions and corollaries in Sec.˜3 of the main paper, with particular focus on demonstrating that in ConvMixer, fixed random impulse spatial filters can achieve comparable performance to learned filters. All experiments adhere to the training protocol proposed in [yoshioka2024visiontransformers], which is specifically tailored for evaluating diverse architectures on small-scale datasets such as CIFAR-10.
To support our theoretical findings in Sec.˜3 concerning the effectiveness of random filters, we train ConvMixer [trockman2022patches] models with an embedding dimension of 256, a depth of 8, and a patch size of 2 on the CIFAR-10 dataset, using spatial filter sizes of 3, 5, and 8. We also include a variant with an embedding dimension of 512 to examine the impact of feature width. We evaluate the end-to-end trained ConvMixer alongside three initialization strategies: random (Corollary 1), impulse (Corollary 2), and box (Corollary 3). Note that in all three cases, only the spatial convolution filters are initialized—the models are evaluated without any training. The box filters use all-one values, effectively performing average pooling. The results are summarized in Tab.˜8.
@@ F.3 Model Convergence Rate
Kernel Size Embedding Dimension == 256 Embedding Dimension == 512 Trained Random Impulse Box Trained Random Impulse Box 3 91.76 90.72 90.68 81.70 92.82 92.15 92.20 81.90 5 92.69 90.87 90.41 80.57 93.90 92.72 91.91 81.19 8 92.34 88.12 87.82 78.95 92.96 90.09 89.61 80.10
Appendix H Attention Maps
Here we provide additional visualization of the attention maps for all 12 layers in Fig.˜7. In particular, our structured initialization method offers different attention peaks on various heads, showing alignment with the impulse structures, while the mimetic initialization only presents main-diagonal peaks, and the default initialization shows little to no patterns. As stated in the main paper, both mimetic and default initialization methods use identical initialization for all attention heads.
Figure 7: Visualization of attention maps in ViT-T using our impulse initialization method, mimetic [trockman2023mimetic], and default [rw2019timm] initializations.